OpenStudy (anonymous):

Somebody please help me, i can do all the other ones except this one x+3y=8 2x-y=2

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

substitution or elimination ? or does it matter ?

OpenStudy (anonymous):

i think substitution @kelliegirl33

OpenStudy (anonymous):

x + 3y = 8 x = -3y + 8 now sub -3y + 8 in for x back into the other equation 2x - y = 2 2(-3y + 8) - y = 2 -6y + 16 - y = 2 -7y + 16 = 2 -7y = 2 - 16 -7y = -14 y = -14/-7 y = 2 now sub 2 in for y in either equation to find x x + 3y = 8 x + 3(2) = 8 x + 6 = 8 x = 8 - 6 x = 2 now we check.. 2x - y = 2 2(2) - 2 = 2 4 - 2 = 2 2 = 2 (correct) so x = 2 and y = 2 questions ?

OpenStudy (anonymous):

Yes why'd you change it from x+3y=8 to x=-3y+8 @kelliegirl33